OpenStudy (anonymous):

find the prime factorization of 11!

OpenStudy (anonymous):

first write it all out, then factor the numbers that are not prime

OpenStudy (anonymous):

wouldn't u factor the numbers that are prime

OpenStudy (anonymous):

you cannot factor a prime number that is why it is called prime

OpenStudy (anonymous):

so ur saying the answer is all of the composite numbers? what?

OpenStudy (anonymous):

start with \[11\times 10\times 9\times 8\times 7\times 6\times 5\times 4\times 3\times 2\]

OpenStudy (anonymous):

then factor the numbers that can be factored, like for example \(10=5\times 2\)

OpenStudy (anonymous):

so it would be 2,3,4,5

OpenStudy (anonymous):

or 2x3x5

OpenStudy (anonymous):

something like the second one

OpenStudy (anonymous):

you have a lot more work to do you have to factor all the numbers that you can

OpenStudy (anonymous):

i think i did...i showed the work on my paper

OpenStudy (anonymous):

\[11!=2^\text{number of even terms +1}\times3^\text{number of terms divisible by 3+1}\\ \qquad\times5^\text{no. of terms divisible by 5}\times7^{1}\times11^{1}\]

OpenStudy (anonymous):

so, \[11!=2^8\times3^4\times5^2\times7\times11\]

OpenStudy (anonymous):

thanks, i got it